Order: Integrilin 180 mcg/kg IV bolus initially. Infuse over 2 minutes. client weighs 154 lb. Available: 2mg/ml. How many ml of the IV bolus is needed to infuse?
Answer #1

Order Integrilin 180mcg/kg

Here weight in pound ie, 154lb

Weight in kilogram 154/2.2=70

So the dose of Integrilin =180×70=12600mcg

One microgram = 0.001milligram

So that, 12600/1000=12.6gm

Here available dose=2mg/ml

2mg=1ml

12.6mg=? ml

Cross multiply, 12.6/2=6.3ml or

There is an equation Required dose/Available dose× Amount to be dissolved in available dose

Ie, 12.6/2×1=6.3ml